The formula gives the height h in feet of an object after t seconds. Here is the initial velocity of the object expressed in feet per second. Suppose a toy rocket was launched from the ground straight up with an initial velocity of 64 ft/sec. How many seconds after launch was the rocket 48 feet above the ground?
Answer by stanbon(75887)

The formula gives the height h in feet of an object after t seconds. Here is the initial velocity of the object expressed in feet per second. Suppose a toy rocket was launched from the ground straight up with an initial velocity of 64 ft/sec. How many seconds after launch was the rocket 48 feet above the ground?
------------
h(t) = -16t^2 + 64t
---------------------
-16t^2 + 64t = 48
-16t^2 + 64t - 48 = 0
Divide thru by -16 to get:
t^2 - 4t + 3 = 0
(t-3)(t-1) = 0
t = 1 sec or t = 3 seconds
--------------------------------
At 1 sec and at 3 seconds theobjectwas 48 ft above the ground.
--------------------------------
Cheers,
Stan H.